OpenStudy (anonymous):

A boy of 50 kg jumps off a boat (30 kg), causing the boat to move to the right at 2.0 m/s. In what direction and at what velocity does the boy move? (Remember to include - to indicate direction if necessary.)

OpenStudy (anonymous):

30 x 2)/50 = 1.2m/sec., the boy moves. I hate "conventions", but the boat moves in the opposite direction to the boy. So you can apply + or - to it, to suit whichever direction you may think is "positive".

OpenStudy (shamim):

Use m1*u1+m2*u2=m1*v1+m2*v2

OpenStudy (shamim):

Given m1=50kg m2=30g u1=0 u2=0 v2=-2m/s v1=?

OpenStudy (shamim):

Response plz!!!